# A SSM is polymerized from multivariate time series

### v1

@@ Abstract
For multivariate time series (MTS) tasks, previous state space models (SSMs) followed the modeling paradigm of Transformer-based methods. However, none of them explicitly model the complex dependencies of MTS: the Channel Dependency variations with Time (CDT). In view of this, we delve into the derivation of SSM, which involves approximating continuously updated functions by orthogonal function bases. We then develop Poly-Mamba, a novel method for MTS forecasting. Its core concept is to expand the original orthogonal function basis space into a multivariate orthogonal function space containing variable mixing terms, and make a projection on this space so as to explicitly describe the CDT by weighted coefficients. In Poly-Mamba, we propose the Multivariate Orthogonal Polynomial Approximation (MOPA) as a simplified implementation of this concept. For the simple linear relationship between channels, we propose Linear Channel Mixing (LCM) and generate CDT patterns adaptively for different channels through a proposed Order Combining method. Experiments on six real-world datasets demonstrate that Poly-Mamba outperforms the SOTA methods, especially when dealing with datasets having a large number of channels and complex correlations. The codes and log files will be released at: https://github.com/Joeland4/Poly-Mamba.
@@ 1 Introduction
Structured state space models (SSMs) [1] are subquadratic-time foundational architectures compared with Transformers [2], and show great performance with approximately linear complexity on long-range dependency tasks. Previous studies [3] [4] [5] attempted to employ SSM for Multivariate Time Series Forecasting (MTSF), they all follow the Transformer-based MTSF modeling paradigm: learning dependencies between temporal tokens [6] [7] [8] [9], Channel tokens [10] and their concatenation [11]. However, the special complex dependency pattern of MTS is the Channel Dependency variations with Time (CDT), none of these methods explicitly depict it.
@@
  • We develop MOPA, an efficient method based on the idea of extending the orthogonal polynomial function basis space to a multivariate orthogonal polynomial basis space and explicitly characterizing channel dependencies through mapping weights.
@@ 2.1 SSM
The target function $u(t)$, a continuously collected signal, is approximated by a linear combination of standard orthogonal function bases $\left\{g_{n}(t)\right\}_{n=0}^{N}$. For this, HIPPO [12] uses an online function approximation method, that is, mapping each static interval $s\in[a,b]$ to the entire sequence interval $t\in[0,T]$ through the defined Measure. The analytical solution of the combination coefficient $C_{n}$ is:
@@ 4.1 Main Results
Secondly, one of the advantages of the state-space model is the compression of history. When facing longer prediction interval requirements, Poly-Mamba shows more stable prediction results. Moreover, compared with the S-Mamba [4] model which is also based on the state-space but models along the channel direction, Poly-Mamba shows more excellent prediction stability. For example, by comparing the results of prediction lengths $T=\left\{336,720\right\}$ on the Solar and Weather datasets, we can find that Poly-Mamba’s modeling is closer to the complex pattern of them.

### v2

@@ Abstract
For multivariate time series (MTS) tasks, previous state space models (SSMs) followed the modeling paradigm of Transformer-based methods. However, none of them explicitly model the complex dependencies of MTS: the Channel Dependency variations with Time (CDT). In view of this, we delve into the derivation of SSM, which involves approximating continuously updated functions by orthogonal function basis. We then develop Poly-Mamba, a novel method for MTS forecasting. Its core concept is to expand the original orthogonal function basis space into a multivariate orthogonal function space containing variable mixing terms, and make a projection on this space so as to explicitly describe the CDT by weighted coefficients. In Poly-Mamba, we propose the Multivariate Orthogonal Polynomial Approximation (MOPA) as a simplified implementation of this concept. For the simple linear relationship between channels, we propose Linear Channel Mixing (LCM) and generate CDT patterns adaptively for different channels through a proposed Order Combining method. Experiments on six real-world datasets demonstrate that Poly-Mamba outperforms the SOTA methods, especially when dealing with datasets having a large number of channels and complex correlations. The codes and log files will be released at: https://github.com/Joeland4/Poly-Mamba.
@@ 1 Introduction
State space models (SSMs) [1] [15] are subquadratic-time foundational architectures compared with Transformers [2], and shows great performance with approximately linear complexity on long-range dependency tasks. Previous studies [3] [4] [5] attempted to employ SSM for Multivariate Time Series Forecasting (MTSF), they all follow the Transformer-based MTSF modeling paradigm: learning dependencies between temporal tokens [6] [7] [8] [9], Channel tokens [10] and their concatenation [11]. However, the special complex dependency pattern of MTS is the Channel Dependency variations with Time (CDT), none of these methods explicitly depict it.
@@
  • We develop MOPA, an efficient method based on the idea of extending the orthogonal polynomial function basis space to a multivariate orthogonal polynomial basis space and explicitly characterizing channel dependencies with the mapped weights.
@@ 2.1 SSM
The target function $u(t)$, a continuously collected signal, is approximated by a linear combination of standard orthogonal function basis $\left\{g_{n}(t)\right\}_{n=0}^{N}$. For this, HIPPO [12] uses an online function approximation method, that is, mapping each static interval $s\in[a,b]$ to the entire sequence interval $t\in[0,T]$ through the defined Measure. The analytical solution of the combination coefficient $C_{n}$ is:
@@ 4.1 Main Results
Secondly, one of the advantages of the state space model is the compression of history. When facing longer prediction interval requirements, Poly-Mamba shows more stable prediction results. Moreover, compared with the S-Mamba [4] model which is also based on the state space but models along the channel direction, Poly-Mamba shows more excellent prediction stability. For example, by comparing the results of prediction lengths $T=\left\{336,720\right\}$ on the Solar and Weather datasets, we can find that Poly-Mamba’s modeling is closer to the complex pattern of them.